ganeshie8 (ganeshie8):

@ikram002p \(\large 13x \equiv 1 (\mod{16})\) how to find the inverse quick without using euclid Answer : 5

OpenStudy (ikram002p):

ok lets try numbers from (1_15) untill we got 1 or -1

ganeshie8 (ganeshie8):

13(1) = -3 (mod 16)

ganeshie8 (ganeshie8):

-1 is useful is it cuz we can square it then to make it 1 hmm

OpenStudy (ikram002p):

13(5) =1 mod 16

ganeshie8 (ganeshie8):

13(2) = 10 (mod 16)

ganeshie8 (ganeshie8):

13(3) = 7 (mod 16)

ganeshie8 (ganeshie8):

13(4) = 4 (mod 16)

ganeshie8 (ganeshie8):

we need to keep guessing like this is it ?

OpenStudy (ikram002p):

will in this gussing method its usfull oly if u got 1 , -1 when numbers is small like this .. but for large num ecliud is much bettter

ganeshie8 (ganeshie8):

oh ok nice :)

ganeshie8 (ganeshie8):

thank you !

OpenStudy (ikram002p):

wlc :)